# On-the-Fly Adaptation to Quantization: Configuration-Aware LoRA for Efficient Fine-Tuning of Quantized LLMs

### v1

@@ 5.2 Research Observations and Experiments
Table 2 demonstrates that CoA-LoRA delivers strong accuracy while being more time-efficient than one-to-one fine-tuning methods (Q-LoRA and LQ-LoRA). This is due to its joint optimization of quantization configurations and configuration-aware model across multiple settings, which induces mutually reinforcing improvements across configurations rather than treating each configuration independently. Concretely, CoA-LoRA learns the ability to adjust LoRA adapters to all considered configurations in roughly one hour, whereas Q-LoRA and LQ-LoRA require 20–40 minutes of fine-tuning per configuration, so their total fine-tuning time increases proportionally with the number of configurations. Shared-LoRA reduces the total fine-tuning time by training one shared set of adapters across multiple quantized settings, but this efficiency comes with degraded accuracy (for example, accuracy gaps of $15.93\%$ on QNLI and $6.14\%$ on SST-2). Taken together, these results show that CoA-LoRA attains the favorable combination of (I) competitive or superior accuracy relative to state-of-the-art methods, and (II) substantially lower fine-tuning time compared to one-to-one fine-tuning methods, while avoiding the pronounced accuracy instability seen in Shared-LoRA.
@@
To assess scalability, Fig. 6 presents results on three pretrained models ranging from 1.5B to 7B parameters, with $N=196,252,224$ layers, respectively. Across all sizes, we observe consistent trends: CoA-LoRA not only outperforms Shared-LoRA but also achieves accuracy comparable to or surpassing other state-of-the-art methods. Notably, GPTQ-LoRA shows curves close to CoA-LoRA on Qwen2.5-1.5B and Qwen2.5-3B, but this requires one-to-one fine-tuning at each bit-width. As model size and the number of configurations grow, this cost scales linearly, whereas CoA-LoRA achieves strong results with a single training process, demonstrating its scalability to larger LLMs.
@@
Figure 8: Performance comparison with and without configuration search across four tasks.
@@ 5.3 Ablation and Sensitivity Analysis
Effect of Configuration Search. Fig. 8 illustrates the impact of applying Gaussian process–based optimization to the configuration set. We observe that this optimization substantially enhances the ability of the configuration-aware model to adjust low-rank matrices, with particularly notable gains on QNLI and SST-2 tasks, where it yields improvements of nearly $2\%$. On MNLI and QQP, the improvements are smaller but still consistent, reflecting that while all tasks benefit from optimization, the extent of improvement varies across datasets.
Performance under Different Ranks. Table 3 reports the hypervolume (HV) and accuracy gap across different ranks. CoA-LoRA consistently achieves the best HV values, indicating that it scales well with the output dimension of the configuration-aware model. In terms of accuracy gap, Shared-LoRA underperforms Q-LoRA in 7 out of 12 cases, whereas CoA-LoRA surpasses Q-LoRA in 11 out of 12 cases. Importantly, Q-LoRA and LQ-LoRA require a separate fine-tuning run for each configuration, while CoA-LoRA reaches comparable or superior results with only a single training process. This advantage comes from both optimizing quantization configurations and training across multiple settings, producing synergistic gains.
Table 3: Hypervolume (HV) and accuracy gap measured at ranks 32, 64, and 128 for four tasks. The best-performing value for each task is highlighted in bold. The accuracy gap is computed relative to Q-LoRA, and the dash (“-”) indicates the baseline itself.
@@ 6 Conclusion
In this work, we present CoA-LoRA, a configuration-aware approach that enables on-the-fly adjustment of low-rank adapters for arbitrary quantization configurations. Experiments show that CoA-LoRA consistently outperforms state-of-the-art methods, achieving accuracy gains of $2.36\%–10.37\%$ over Q-LoRA across four GLUE tasks, and HV improvements of $3\%–8\%$ compared to LQ-LoRA. Importantly, CoA-LoRA allows real-time adaptation of low-rank adapters to arbitrary configurations without additional fine-tuning, while maintaining stable performance across tasks and strong generalization to unseen configurations. These results demonstrate CoA-LoRA’s efficiency and robustness for deploying large-scale LLMs under heterogeneous device capabilities.
@@ Appendix A Algorithm Details
A.1 Generation of the Initial Quantization Configuration Set
@@
s.t. $\texttt{storage}(\widetilde{\mathbf{W}}_{\mathbf{X}})\leq b,$
@@
where $\texttt{storage}(\widetilde{\mathbf{W}}_{\mathbf{X}})$ denotes the memory cost (in bits) required to store the quantized weights $\widetilde{\mathbf{W}}_{\mathbf{X}}$. Intuitively, this procedure seeks a binary quantization assignment $\mathbf{X}$ that balances quantization (via the storage constraint) and reconstruction error (via the low-rank residual), providing a well-initialized configuration for subsequent LoRA fine-tuning.
@@
A.2 Quantization Configuration Selection for a Given Bit-Width
@@
s.t. $\texttt{storage}(\mathbf{C})=b,$
where $\mathrm{dist}(\mathbf{C},\mathcal{C})=\min_{\mathbf{C}^{\prime}\in\mathcal{C}}\|\mathbf{C}-\mathbf{C}^{\prime}\|_{F}$ measures the distance to the nearest configuration in $\mathcal{C}$, and $\texttt{storage}(\mathbf{C})$ denotes the overall average bit-width of $\mathbf{C}$.
Intuitively, this procedure leverages the representativeness of the final configuration set: it ensures that the selected configuration is both feasible for the target bit-width and similar to the training configurations, thereby ensuring that the configuration performs well under the specified bit-width. As shown in Fig. A.1, this procedure ensures that the configuration-aware model produces LoRA adjustments that generalize effectively to new configurations with nearby bit-widths.
@@ 10   Sample $\mathbf{C}\sim\mathcal{C}$;
Figure C.1: Correlations of two types of low-rank matrices after LoRA fine-tuning on RoBERTa-Large, measured separately for models quantized to 2, 3, 4, 5, and 6 bits.
@@ Appendix B Experimental Details
For the computation of Hypervolume (HV) (Zitzler & Thiele, 1999), we first collect the performance metrics for each algorithm. The second objective $f_{2}$ can be either accuracy or perplexity depending on the experiment.
@@ C.1 Similarity of Low-Rank Matrices Across Different Bit-Width Models
Fig. C.1 presents the block-wise averaged similarity between $\mathbf{L}_{1}$ and $\mathbf{L}_{2}$ after quantizing RoBERTa-Large to 2, 3, 4, 5, and 6 bits using Eq. 11, followed by low-rank fine-tuning. We observe that the shared components across different configurations are mostly captured in $\mathbf{L}_{2}$, while the configuration-specific knowledge is primarily encoded in $\mathbf{L}_{1}$. This observation motivates the design of the configuration-aware model $\bm{\theta}$, where the model outputs an $r\times r$ matrix $\mathbf{U}_{\bm{\theta}}$ to directly transform $\mathbf{L}_{2}$ into $\mathbf{U}_{\bm{\theta}}\mathbf{L}_{2}$.
Table C.2: Comparison of hypervolume (HV) and average decrease in perplexity (lower is better) relative to Q-LoRA across three LLMs.
@@ C.4 Visualization Comparison of Methods under Varying Ranks on RoBERTa-Large
Fig. C.3 visualizes the Pareto fronts of CoA-LoRA and four baselines, showing that CoA-LoRA achieves accuracy comparable to, and often surpassing, Q-LoRA and LQ-LoRA, which rely on one-to-one fine-tuning after quantization.

### v2

@@ 5.2 Research Observations and Experiments
Table 2 demonstrates that CoA-LoRA delivers strong accuracy while being more time-efficient than one-to-one fine-tuning methods (QLoRA and LQ-LoRA). This is due to its joint optimization of quantization configurations and configuration-aware model across multiple settings, which induces mutually reinforcing improvements across configurations rather than treating each configuration independently. Concretely, CoA-LoRA learns the ability to adjust LoRA adapters to all considered configurations in roughly one hour, whereas QLoRA and LQ-LoRA require 20–40 minutes of fine-tuning per configuration, so their total fine-tuning time increases proportionally with the number of configurations. Shared-LoRA reduces the total fine-tuning time by training one shared set of adapters across multiple quantized settings, but this efficiency comes with degraded accuracy (for example, accuracy gaps of $5.06\%$ on SST-2 and $1.18\%$ on QQP). Fig. 5 shows that CoA-LoRA maintains leading performance across various average bit-widths. Taken together, these results show that CoA-LoRA attains the favorable combination of (I) competitive or superior accuracy relative to state-of-the-art methods, and (II) substantially lower fine-tuning time compared to one-to-one fine-tuning methods, while avoiding the pronounced accuracy instability seen in Shared-LoRA.
@@
To assess scalability, Fig. 6 presents results on three pretrained models ranging from 1.5B to 7B parameters, with $N=196,252,224$ layers, respectively. Across all sizes, we observe consistent trends: CoA-LoRA not only outperforms Shared-LoRA but also achieves accuracy comparable to or surpassing other state-of-the-art methods. Notably, GPTQLoRA shows curves close to CoA-LoRA on Qwen2.5-1.5B and Qwen2.5-3B, but this requires one-to-one fine-tuning at each bit-width. As model size and the number of configurations grow, this cost scales linearly, whereas CoA-LoRA achieves strong results with a single training process, demonstrating its scalability to larger LLMs.
@@
Figure 8: Effect of configuration search in CoA-LoRA.
@@ 5.3 Ablation and Sensitivity Analysis
Effect of Configuration Search. Fig. 8 illustrates the impact of applying Gaussian process–based optimization to the configuration set. To enable a more informative comparison, we further introduce a random extension baseline, where each epoch randomly adds ten new quantization configurations to the set. This comparison leads to two key findings. (I) Configuration search enhances the ability of the configuration-aware model to adjust low-rank matrices. Compared with the no–configuration-search setting, we observe substantial gains on QNLI and SST-2, reaching nearly $2\%$. Improvements on MNLI and QQP are smaller but remain consistent, indicating that all tasks benefit from optimization, though to varying degrees. (II) Configuration search supplies high-quality candidate configurations that effectively strengthen the configuration-aware model. In contrast, random extension achieves performance comparable to the no–configuration-search baseline but remains substantially below that of configuration search. This outcome suggests that increasing the configuration set with arbitrary configurations is ineffective—performance gains depend on introducing high-quality candidates obtained via guided optimization (i.e., Pareto-based configuration search).
Performance under Different Ranks. Table 3 reports the hypervolume (HV) and accuracy gap across different ranks. CoA-LoRA consistently achieves the best HV values, indicating that it scales well with the output dimension of the configuration-aware model. In terms of accuracy gap, Shared-LoRA underperforms QLoRA in 7 out of 12 cases, whereas CoA-LoRA surpasses QLoRA in 11 out of 12 cases. Importantly, QLoRA and LQ-LoRA require a separate fine-tuning run for each configuration, while CoA-LoRA reaches comparable or superior results with only a single training process. This advantage comes from both optimizing quantization configurations and training across multiple settings, producing synergistic gains.
Table 3: Hypervolume (HV) and accuracy gap measured at ranks 32, 64, and 128 for four tasks. The best-performing value is highlighted in bold. The dash (“-”) indicates the baseline itself.
@@
Comparison under Integer Mixed-Precision Quantization. Table 4 reports results for integer mixed-precision quantization. CoA-LoRA consistently achieves strong performance across the four tasks, attaining the best results in 10 out of 16 task–bit combinations. Notably, it maintains competitive accuracy even at lower bit-widths (e.g., 3 bit), whereas other methods such as QLoRA and LQLoRA tend to degrade under aggressive quantization. Across tasks, CoA-LoRA also exhibits more stable improvements as bit-width increases, while Shared-LoRA and QLoRA show more variability between datasets. These observations suggest that CoA-LoRA’s robustness and versatility, demonstrating that its advantages extend beyond floating-point settings to integer-based settings.
Table 4: Accuracy comparison under integer mixed-precision quantization (per-layer choices: int2, int3, int4, int8). Columns correspond to methods, and values within each cell are reported in the order of tasks: QNLI / MNLI / SST-2 / QQP.
Avg. Bit QLoRA LQ-LoRA Shared-LoRA CoA-LoRA
3 0.8537 / 0.8153 / 0.8291 / 0.7353 0.8552 / 0.7397 / 0.9208 / 0.7777 0.7772 / 0.6618 / 0.8199 / 0.7311 0.8616 / 0.8099 / 0.8589 / 0.7621
4 0.8828 / 0.8491 / 0.9116 / 0.7469 0.8762 / 0.7913 / 0.9185 / 0.7952 0.8762 / 0.7632 / 0.8486 / 0.7626 0.8841 / 0.8441 / 0.9254 / 0.8223
5 0.8859 / 0.8554 / 0.9105 / 0.7501 0.8761 / 0.8481 / 0.9369 / 0.8000 0.8777 / 0.7674 / 0.8383 / 0.7403 0.8863 / 0.8533 / 0.9346 / 0.8160
6 0.8833 / 0.8466 / 0.9174 / 0.7441 0.8744 / 0.8484 / 0.9323 / 0.7918 0.8790 / 0.7704 / 0.8440 / 0.7443 0.8925 / 0.8558 / 0.9323 / 0.8214
@@ 6 Conclusion
In this work, we present CoA-LoRA, a configuration-aware approach that enables on-the-fly adjustment of low-rank adapters for arbitrary quantization configurations. Experiments show that CoA-LoRA consistently outperforms state-of-the-art methods, achieving accuracy gains of $1.74\%–8.89\%$ over QLoRA across four GLUE tasks, and HV improvements of $2\%–7\%$ compared to LQ-LoRA. Importantly, CoA-LoRA allows real-time adaptation of low-rank adapters to arbitrary configurations without additional fine-tuning, while maintaining stable performance across tasks and strong generalization to unseen configurations. These results demonstrate CoA-LoRA’s efficiency and robustness for deploying large-scale LLMs under heterogeneous device capabilities.
Acknowledgments
This work was supported in part by the Guangdong Basic and Applied Basic Research Foundation under Grant 2023A1515012819 and the National Natural Science Foundation of China under Grant 62202214, and in part by the UGC General Research Fund no. 17209822 and the Innovation and Technology Commission Fund no. ITS/383/23FP from Hong Kong.
@@ Appendix A Algorithm Details
A.1 Configuration-Aware Mapper Architecture
The configuration-aware mapper is a conditional network that generates layer-specific adjustment matrices for LoRA parameters. The network consists of separate embeddings encoding layer and quantization information, followed by a two-layer MLP that produces a square adjustment matrix for each layer. The architecture details are summarized below:
(i) Input Embeddings. For each layer, several types of embeddings are constructed: Module type: For RoBERTa-Large: query,key,value,output.dense,intermediate.densequery,key,value,output.dense,intermediate.dense; for LLaMA and Qwen2.5: $q_{proj},k_{proj},v_{proj},o_{proj},gate_{proj},down_{proj},up_{proj}$. Quantization parameters: $b_{0}$, $b_{1}$, $b_{2}$ (e.g., fp16, bf16, fp32), $B_{0}$, $B_{1}$. Block position: the block index.
(ii) Embedding Dimension. Each embedding has a dimensionality of $d=8$.
(iii) MLP Mapping. The concatenated embeddings form the input to a two-layer MLP with LayerNorm and SiLU activation. The MLP has an input dimension of $7\cdot d$, where the factor 7 corresponds to the concatenation of seven embeddings: one for the module type, five for quantization parameters, and one for the block position. The MLP has a hidden dimension of 128 and an output dimension of $r^{2}$, which is reshaped into an $r\times r$ adjustment matrix. The output is reshaped into an $r\times r$ adjustment matrix.
(iv) Output Adjustment. The adjustment matrix is scaled by a learnable parameter and added to the identity matrix to stabilize the transformation.
The base LoRA adapters are constructed by performing SVD on the difference between the full-precision and quantized weights. Further details can be found in Appendix A.2.
Adjustment matrices are generated independently for each layer, enabling on-the-fly and parallel computation for all layers. We list the trainable parameters of the configuration-aware mapper in Table A.1. The results show that the configuration-aware mapper accounts for less than 0.5% of the total parameters.
Table A.1: Parameter breakdown for different models: total model parameters / LoRA parameters (rank = 128) / configuration-aware mapper parameters.
Model Parameters Percentage
RoBERTa-Large 414,934,577 / 60,845,617 / 2,121,265 100% / 14% / 0.5%
Qwen2.5-1.5B 1,693,553,729 / 149,839,425 / 2,121,281 100% / 8% / 0.1%
Qwen2.5-3B 3,327,528,513 / 241,589,825 / 2,121,281 100% / 7% / 0.06%
Llama-2-7B 7,060,352,577 / 321,936,961 / 2,121,281 100% / 4% / 0.03%
A.2 Generation of the Initial Quantization Configuration Set
@@
s.t. $f_{2}(\widetilde{\mathbf{W}}_{\mathbf{X}})\leq\text{budget},$
@@
where $f_{2}(\widetilde{\mathbf{W}}_{\mathbf{X}})$ denotes average bit of $\widetilde{\mathbf{W}}_{\mathbf{X}}$. It can be computed as
$f_{2}(\mathbf{C})=\frac{\sum_{i=1}^{L}(b_{0}+\frac{b_{1}}{B_{0}}+\frac{b_{2}}{B_{0}B_{1}})\cdot w_{i}}{\sum_{i=1}^{L}w_{i}},$ (12)
where $w_{i}$ represents the number of parameters in layer $i$ and $\mathbf{C}=[b_{0},b_{1},b_{2},B_{0},B_{1}]$. Here, $b_{2}$ denotes the bitwidth of the precision type. If the type is fp16 or bf16, we set $b_{2}=16$; if it is fp32, we set $b_{2}=32$. Intuitively, this procedure seeks a binary quantization assignment $\mathbf{X}$ that balances quantization (via the storage constraint) and reconstruction error (via the low-rank residual), providing a well-initialized configuration for subsequent LoRA fine-tuning.
@@
A.3 Quantization Configuration Selection for a Given Bit-Width
@@
s.t. $f_{2}(\mathbf{C})=b,$
where $\mathrm{dist}(\mathbf{C},\mathcal{C})=\min_{\mathbf{C}^{\prime}\in\mathcal{C}}\|\mathbf{C}-\mathbf{C}^{\prime}\|_{F}$ measures the distance to the nearest configuration in $\mathcal{C}$.
Intuitively, problem (13) selects the representativeness of the final configuration set: it ensures that the selected configuration is both feasible for the target bit-width and similar to the training configurations, thereby ensuring that the configuration performs well under the specified bit-width. As shown in Fig. A.1, this procedure ensures that the configuration-aware model produces LoRA adjustments that generalize effectively to new configurations with nearby bit-widths.
@@ A.4 Training Procedure
Figure C.1: Correlations of two types of low-rank matrices after LoRA fine-tuning on RoBERTa-Large for models quantized to 2, 3, 4, 5, and 6 bits, measured on the QNLI dataset. The three panels correspond to rank 32 (top), rank 64 (middle), and rank 128 (middle). Each panel shows two rows representing the block-wise averaged similarity for $\mathbf{L}_{1}$ and $\mathbf{L}_{2}$.
Figure C.2: Correlations of two types of low-rank matrices after LoRA fine-tuning for models quantized to 2, 3, 4, 5, and 6 bits, measured on the C4 dataset. The three panels correspond to Qwen-1.5B (top), Qwen-3B (middle), and Llama-2-7B (bottom). Each panel shows two rows representing the block-wise averaged similarity for $\mathbf{L}_{1}$ and $\mathbf{L}_{2}$.
@@ Appendix B Experimental Details
For the computation of Hypervolume (HV) (Zitzler and Thiele, 1999), we first collect the performance metrics for each algorithm. The second objective $f_{2}$ can be either accuracy or perplexity depending on the experiment.
@@ C.1 Similarity of Low-Rank Matrices Across Different Bit-Width Models
We evaluated the correlations between $\mathbf{L}_{1}$ and $\mathbf{L}_{2}$ under different ranks and across multiple models. Figures C.1 and C.2 show the block-wise averaged similarity between $\mathbf{L}_{1}$ and $\mathbf{L}_{2}$ after quantizing RoBERTa-Large to 2, 3, 4, 5, and 6 bits using Eq. 11, followed by low-rank fine-tuning. We observe that the shared components across different configurations are mostly captured in $\mathbf{L}_{2}$, while the configuration-specific knowledge is primarily encoded in $\mathbf{L}_{1}$. This observation motivates the design of the configuration-aware model $\bm{\theta}$, where the model outputs an $r\times r$ matrix $\mathbf{U}_{\bm{\theta}}$ to directly transform $\mathbf{L}_{2}$ into $\mathbf{U}_{\bm{\theta}}\mathbf{L}_{2}$.
Table C.2: Comparison of hypervolume (HV) and average decrease in perplexity (lower is better) relative to QLoRA across three LLMs.
@@ C.4 Visualization Comparison of Methods under Varying Ranks on RoBERTa-Large
Fig. C.4 visualizes the Pareto fronts of CoA-LoRA and four baselines, showing that CoA-LoRA achieves accuracy comparable to, and often surpassing, QLoRA and LQ-LoRA, which rely on one-to-one fine-tuning after quantization.
C.5 Comparison under Llama-2-13B
We evaluated CoA-LoRA on LLaMA-2-13B across multiple quantization bits. Table C.4 reports the perplexity for different methods. CoA-LoRA achieves performance comparable to one-to-one full fine-tuning across most bit settings, including low-bit configurations (2.5–3.5 bits). These results indicate that CoA-LoRA can scale to larger models while maintaining stable and efficient adaptation to different quantization levels.
Table C.4: Perplexity comparison under different quantization bits (2.5, 3, 3.5, 4) on LLaMA-2-13B with rank 64. ”-” denotes that GPTQ does not officially support 2.5-bit and 3.5-bit quantization.
Avg. Bit QLoRA LQ-LoRA GPTQ-LoRA Shared-LoRA CoA-LoRA
2.5 9.11 9.23 - 8.87 8.65
3 7.48 8.00 7.95 8.18 7.58
3.5 7.09 7.52 - 7.59 7.15
4 6.89 7.32 7.15 7.28 6.99
C.6 Comparison of Llama-2-7B under different ranks
Table C.5 reports the performance of different methods across various ranks and bit-widths. CoA-LoRA consistently exhibits stable results across all bit settings. At rank 32, its performance remains steady, while Shared-LoRA shows extreme fluctuations at bit 2.5. Similar patterns are observed at ranks 64 and 128, where CoA-LoRA avoids the large variations exhibited by LQ-LoRA and Shared-LoRA under low bit conditions. These results indicate that CoA-LoRA preserves performance under aggressive quantization and delivers reliable behavior across different Ranks.
Table C.5: Performance of LoRA quantization methods on LLaMA-2-7B across different ranks (32, 64, 128) and bit-widths (2.5–4). ”-” denotes that GPTQ does not officially support 2.5-bit and 3.5-bit quantization.
Rank Avg. Bit QLoRA LQ-LoRA GPTQ-LoRA Shared-LoRA CoA-LoRA
32 2.5 9.48 9.94 - 33.26 8.85
3 8.28 8.63 8.63 8.43 8.76
3.5 7.92 7.59 - 8.01 7.56
4 7.70 7.37 7.73 7.91 7.49
64 2.5 11.02 21.29 - 8.87 8.48
3 7.67 8.28 8.62 8.01 7.91
3.5 7.27 8.26 - 7.52 7.52
4 7.06 8.09 7.72 7.27 7.23
128 2.5 9.46 9.79 - 24.93 9.21
3 7.64 7.86 8.66 8.78 7.77
3.5 7.91 7.51 - 7.67 7.51
4 7.71 7.33 7.72 7.54 7.39
Figure C.5: Configuration set sizes of CoA-LoRA across four tasks.